OpenStudy (anonymous):

Which set of coordinates is a solution of the inequality, y ≥ 5?

OpenStudy (anonymous):

A. (-1, 5) B. (3, -2) C. (4, 3) D. (7, 0)

OpenStudy (anonymous):

Is it \(y \ge 5\) or is there an 'x' in there somewhere?

OpenStudy (anonymous):

If there is no 'x', look at the 2nd number in each option and see which one is greater than or equal to 5.

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

No, 3 isn't greater than or equal to 5.

OpenStudy (anonymous):

a?

OpenStudy (anonymous):

im confuseed

OpenStudy (anonymous):

Yep, A is correct. (-1, 5) ^ This is 'y', so we plug that in and see if it's true: \(y \ge 5\) \(5 \ge 5\) 5 is equal to 5, so it is correct.